# Collaborative Representation Learning for Alignment of Tactile, Language, and Vision Modalities

### v1

@@ 3.4 Theoretical Guarantees
We analyze TLV-CoRe’s properties under standard optimization assumptions (patel2022global; xu2025psmgd; lei2019stochastic; pham2020proxsarah). Let $\mathcal{L}(\Theta)$ denote the training objective. We make the following assumptions to facilitate analysis:
@@
Assumption 3.2 (Polyak–Łojasiewicz (PL) Condition (karimi2016linear)).
@@
TacQuad GelSight, DIGIT, DuraGel, GelSight Mini 55k
@@
Table 3: Performance (%) comparison of different methods under three evaluation protocols: intra-sensor evaluation, cross-sensor generalization, and multi-sensor generalization. $\dagger$} Note that we follow the default configuration of the TLV-Link repository, applying LoRA (hu2022lora) to fine-tune the tactile and vision encoders while keeping the language encoder frozen. For methods that cannot handle missing modalities, the tactile modality is used as a substitute for the missing vision modality in the Octopi dataset.
@@ 4 Experiments
Under the RSS evaluation framework, our experiments primarily focus on tactile representations and adopt a comprehensive and consistent linear probing approach (cheng2024touch100k; tacquad; lei2024vitlen) to evaluate the quality of tactile representations from different methods across three evaluation protocols. The specific evaluation protocols include intra-sensor evaluation, cross-sensor generalization, and multi-sensor generalization, across two tasks—material property identification and robot grasping prediction.
@@ 5.1 Real-World Tactile Datasets
We train various tactile representation learning methods on single-sensor datasets (Touch and Go (TAG) (yang2022touch), SSVTP (kerr2022ssvtp), TVL (fu2024tvl), and Octopi (yu2024octopi)) and a multi-sensor dataset (TacQuad (tacquad)). For material property identification, we select three test subsets from TAG, along with ObjectFolder 1.0 (OF 1.0) (gao2022objectfolder1) and ObjectFolder 2.0 (OF 2.0) (gao2022objectfolder2), as downstream evaluation datasets. TAG includes three tactile classification tasks: material (20 classes), roughness (2 classes), and hardness (2 classes). OF 1.0 and OF 2.0 focus on material classification (7 classes). The two datasets splits follow the setup in (yang2024binding). For robot grasping prediction, we evaluate on The Feeling of Success (Feel) dataset (calandra2017feeling). Following (tacquad), we use frames from the left and right tactile sensors during the grasping process as input to predict whether a grasp is successful. Based on (yang2024binding; cheng2024touch100k), we split the Feel dataset by object into training, validation, and test sets with a ratio of 8:1:1. Table 2 presents the statistics of the aforementioned datasets. Notably, the language modality in TAG is generated by GPT-4o, and textual descriptions in TVL, SSVTP, and Octopi are extended, as detailed in (tacquad).
@@ 5.2 Tactile-Language-Vision Baselines
We compare the proposed TLV-CoRe method with state-of-the-art multimodal tactile approaches, including single-sensor methods TLV-Link (cheng2024touch100k) and VIT-LENS-2 (lei2024vitlen), as well as the multi-sensor method AnyTouch (tacquad). Notably, all these methods use OpenCLIP-large (cherti2023reproducible) to initialize the tactile, language, and vision encoders, which allows for a fair comparison among different multimodal tactile approaches without the confounding factor of varying base models. Specifically, both the tactile and vision encoders adopt a 24-layer, 1024-dimensional Vision Transformer (ViT) (dosovitskiy2020image) with a patch size of 14, while the language encoder uses a 12-layer, 768-dimensional Transformer architecture (vaswani2017attention). It’s noted that since UniTouch (yang2024binding) is not fully open-source, we do not use it as a baseline in our work.
@@ 5.3 Implementation Details
We train various tactile models for 12 epochs, except for VIT-LENS-2 which is trained for 80 epochs to ensure convergence, using the AdamW optimizer (loshchilov2017decoupled) with an initial learning rate of 2e-4 and momentum parameters $\beta_{1},\beta_{2}=0.9,0.98$ on two NVIDIA A800 GPUs. The first stage of AnyTouch is trained for 20 epochs. All models are trained with linear probing for 50 epochs. For both TLV-Link and AnyTouch, the LoRA rank is set to 16. VIT-LENS-2 adopts the VIT-LENSL architecture. For the proposed TLV-CoRe model, the shared projection layer dimension is set to 32. We apply UBA to all layers of the language encoder and introduce UBA to the tactile and vision encoders starting from layer $k=13$, ensuring one-to-one correspondence across the three modalities. In cross-sensor tactile decoupling, the hyperparameter $\lambda$ is set to 0.1. Unless otherwise specified, all models are trained with a batch size of 64, and other hyperparameters follow the original papers. Following previous work (yang2022touch; yang2024binding; cheng2024touch100k; tacquad; lei2024vitlen), we use accuracy as the evaluation metric for classification tasks.
@@ 5.5 Tactile $\times$ Vision $\Rightarrow$ Synergy Analysis
The goal of cross-modal alignment is to achieve information complementarity between modalities, rather than sacrificing the representation quality of each modality (wang2022deep; dufumier2024align). In other words, the aim is to enable synergy across modalities. To evaluate how well different methods achieve this synergy between tactile and vision modalities, we design a modal cross-evaluation. We select models trained on the TAG, Octopi, and TacQuad datasets in Sec. 5.4 as the evaluation subjects. For tactile tasks, we choose three subtasks from the TAG dataset and evaluate them using the corresponding models’ vision encoders. For vision tasks, we use three image classification datasets—CIFAR-10, CIFAR-100 (krizhevsky2009learning), and ImgNetDogs (dog-breed-identification)—and evaluate them with the tactile encoders. Table 4 shows the following observations: (1) In tactile tasks, vision encoders generally perform worse than tactile encoders, which is expected given their modality focus. (2) In vision tasks, most methods demonstrate limited performance, possibly due to the setting of only 50 linear probing training epochs, which may indicate that a model with only a learnable classification head may not have sufficiently learned the new visual information. (3) Notably, the proposed TLV-CoRe achieves the best performance across both types of tasks. We attribute this to the introduction of the UBA module, which bridges the modality branches via parameter sharing. This facilitates a unified feature space, enabling smooth gradient flow and efficient information transfer and complementarity across modalities, while also accelerating convergence. These results further validate the analysis in Sec. 3.4.
@@ 5.6 Different Batch Sizes $\Rightarrow$ Stability Analysis
In CLIP-based contrastive methods, different batch sizes lead to significant variations in the number of negative samples. As illustrated in Fig.3, batch size significantly affects the performance of various methods trained by TacQuad dataset, evaluated across three subsets of TAG dataset. Notably, larger batch sizes tend to yield better performance, which aligns with observations from previous studies (chen2020simple; kerr2022self). However, we also observe that when the batch size reaches 128, the performance of TLV-Link and VIT-LENS-2 no longer improves and may even decline. This is because a larger batch contains more data from different sensors, making it harder for single-sensor methods to handle the increased heterogeneity in tactile data. Overall, the proposed TLV-CoRe exhibits a smoother performance curve and demonstrates superior stability. This observation is consistent with the analysis presented in Sec. 3.4.
@@ 5.7 Ablation Study
We perform an ablation study on the components of TLV-CoRe trained by TacQuad. As shown in Fig. 4, removing the SAM leads to a performance drop, and eliminating the decoupled loss $\mathcal{L}_{DL}$ causes an even greater decline, highlighting the crucial role of decoupled learning in handling multi-sensor data. Additionally, removing the UBA module results in the worst performance, underscoring its critical role in the overall architecture.
Additionally, we further investigate the impact of hyperparameters on TLV-CoRe trained by TacQuad dataset: the decoupled coefficient $\lambda_{DL}$, the number of UBA $N_{UBA}$, and the dimension of UBA shared layer $d_{shared}$. As shown in Table 5, both $\lambda_{DL}$ and $d_{shared}$ involve a trade-off between performance and generalization: a low $\lambda_{DL}$ weakens the decoupled effect, while a high value suppresses contrastive learning; a small $d_{shared}$ leads to insufficient sharing, whereas a large one tends to cause overfitting. In contrast, increasing $N_{UBA}$ generally improves performance, which indicates that it’s necessary to fully promote collaboration and communication between different modalities.
@@
To investigate the impact of tactile images with similar styles but from different sensors on model performance, we conduct experiments using the GelSight and DIGIT data from TacQuad dataset, as they exhibit stylistic similarity. Specifically, we randomly sample 5,000 instances from each dataset and additionally sample 2,500 instances from each to form a combined GelSight + DIGIT dataset. Due to the limited data size, we set the batch size to 8. Table 6 compares the performance of the multi-sensor methods AnyTouch and TLV-CoRe. Results show that AnyTouch suffers a performance drop when shifting from single-sensor data (GelSight or DIGIT) to the mixed dataset, indicating its inability to handle style-consistent sensor differences, which leads to degraded multi-sensor representations. In contrast, TLV-CoRe mitigates this issue through decoupled learning and demonstrates more stable performance.
@@ Appendix A Convergence Analysis
To validate our theoretical analysis of convergence properties in Theorem 3.1, we compare the training performance of TLV-CoRe against two state-of-the-art approaches: TLV-Link (cheng2024touch100k) and AnyTouch (tacquad). Fig. 5 shows the test loss trajectories over 12 epochs on the TAG dataset. The results provide strong empirical support for our theoretical claims. First, TLV-CoRe exhibits the fastest convergence rate among all methods, achieving a stable loss value by epoch 8, while other approaches continue to fluctuate. Second, unlike TLV-Link which shows instability after epoch 8 with increasing loss, TLV-CoRe maintains consistent performance. The combined effect of these properties enables TLV-CoRe to achieve both lower final test loss and more reliable optimization behavior, creating a foundation for robust downstream performance in this datasets and others.

### v3

@@ 3.4 Theoretical Guarantees
We analyze TLV-CoRe’s properties under standard optimization assumptions (Patel, Zhang, and Tian 2022; Xu et al. 2025; Lei et al. 2019; Pham et al. 2020). Let $\mathcal{L}(\Theta)$ denote the training objective. We make the following assumptions to facilitate analysis:
@@
Assumption 3.2 (Polyak–Łojasiewicz (PL) Condition (Karimi, Nutini, and Schmidt 2016)).
@@
feng2025anytouch GelSight, DIGIT, DuraGel, GelSight Mini 55k
@@
Table 3: Performance (%) comparison of different methods under three evaluation protocols: intra-sensor evaluation, cross-sensor generalization, and multi-sensor generalization. $\dagger$} Note that we follow the default configuration of the TLV-Link repository, applying LoRA (Hu et al. 2022) to fine-tune the tactile and vision encoders while keeping the language encoder frozen. For methods that cannot handle missing modalities, the tactile modality is used as a substitute for the missing vision modality in the Octopi dataset.
@@ 4 Experiments
Under the RSS evaluation framework, our experiments primarily focus on tactile representations and adopt a comprehensive and consistent linear probing approach (Cheng et al. 2025; Feng et al. 2025; Lei et al. 2024) to evaluate the quality of tactile representations from different methods across three evaluation protocols. The specific evaluation protocols include intra-sensor evaluation, cross-sensor generalization, and multi-sensor generalization, across two tasks—material property identification and robot grasping prediction.
@@ 5.1 Real-World Tactile Datasets
We train various tactile representation learning methods on single-sensor datasets (Touch and Go (TAG) (Yang et al. 2022), SSVTP (Kerr et al. 2022), TVL (Fu et al. 2024), and Octopi (Yu et al. 2024)) and a multi-sensor dataset (feng2025anytouch (Feng et al. 2025)). For material property identification, we select three test subsets from TAG, along with ObjectFolder 1.0 (OF 1.0) (Gao et al. 2021) and ObjectFolder 2.0 (OF 2.0) (Gao et al. 2022), as downstream evaluation datasets. TAG includes three tactile classification tasks: material (20 classes), roughness (2 classes), and hardness (2 classes). OF 1.0 and OF 2.0 focus on material classification (7 classes). The two datasets splits follow the setup in (Yang et al. 2024). For robot grasping prediction, we evaluate on The Feeling of Success (Feel) dataset (Calandra et al. 2017). Following (Feng et al. 2025), we use frames from the left and right tactile sensors during the grasping process as input to predict whether a grasp is successful. Based on (Yang et al. 2024; Cheng et al. 2025), we split the Feel dataset by object into training, validation, and test sets with a ratio of 8:1:1. Table 2 presents the statistics of the aforementioned datasets. Notably, the language modality in TAG is generated by GPT-4o, and textual descriptions in TVL, SSVTP, and Octopi are extended, as detailed in (Feng et al. 2025).
@@ 5.2 Tactile-Language-Vision Baselines
We compare the proposed TLV-CoRe method with state-of-the-art multimodal tactile approaches, including single-sensor methods TLV-Link (Cheng et al. 2025) and VIT-LENS-2 (Lei et al. 2024), as well as the multi-sensor method AnyTouch (Feng et al. 2025). Notably, all these methods use OpenCLIP-large (Cherti et al. 2023) to initialize the tactile, language, and vision encoders, which allows for a fair comparison among different multimodal tactile approaches without the confounding factor of varying base models. Specifically, both the tactile and vision encoders adopt a 24-layer, 1024-dimensional Vision Transformer (ViT) (Dosovitskiy 2020) with a patch size of 14, while the language encoder uses a 12-layer, 768-dimensional Transformer architecture (Vaswani et al. 2017). It’s noted that since UniTouch (Yang et al. 2024) is not fully open-source, we do not use it as a baseline in our work.
@@ 5.3 Implementation Details
We train various tactile models for 12 epochs, except for VIT-LENS-2 which is trained for 80 epochs to ensure convergence, using the AdamW optimizer (Loshchilov and Hutter 2017) with an initial learning rate of 2e-4 and momentum parameters $\beta_{1},\beta_{2}=0.9,0.98$ on two NVIDIA A800 GPUs. The first stage of AnyTouch is trained for 20 epochs. All models are trained with linear probing for 50 epochs. For both TLV-Link and AnyTouch, the LoRA rank is set to 16. VIT-LENS-2 adopts the VIT-LENSL architecture. For the proposed TLV-CoRe model, the shared projection layer dimension is set to 32. We apply UBA to all layers of the language encoder and introduce UBA to the tactile and vision encoders starting from layer $k=13$, ensuring one-to-one correspondence across the three modalities. In cross-sensor tactile decoupling, the hyperparameter $\lambda$ is set to 0.1. Unless otherwise specified, all models are trained with a batch size of 64, and other hyperparameters follow the original papers. Following previous work (Yang et al. 2022, 2024; Cheng et al. 2025; Feng et al. 2025; Lei et al. 2024), we use accuracy as the evaluation metric for classification tasks.
@@ 5.5 Tactile $\times$ Vision $\Rightarrow$ Synergy Analysis
The goal of cross-modal alignment is to achieve information complementarity between modalities, rather than sacrificing the representation quality of each modality (Wang et al. 2022; Dufumier et al. 2024). In other words, the aim is to enable synergy across modalities. To evaluate how well different methods achieve this synergy between tactile and vision modalities, we design a modal cross-evaluation. We select models trained on the TAG, Octopi, and feng2025anytouch datasets in Sec. 5.4 as the evaluation subjects. For tactile tasks, we choose three subtasks from the TAG dataset and evaluate them using the corresponding models’ vision encoders. For vision tasks, we use three image classification datasets—CIFAR-10, CIFAR-100 (Krizhevsky, Hinton et al. 2009), and ImgNetDogs (Ráduly et al. 2018)—and evaluate them with the tactile encoders. Table 4 shows the following observations: (1) In tactile tasks, vision encoders generally perform worse than tactile encoders, which is expected given their modality focus. (2) In vision tasks, most methods demonstrate limited performance, possibly due to the setting of only 50 linear probing training epochs, which may indicate that a model with only a learnable classification head may not have sufficiently learned the new visual information. (3) Notably, the proposed TLV-CoRe achieves the best performance across both types of tasks. We attribute this to the introduction of the UBA module, which bridges the modality branches via parameter sharing. This facilitates a unified feature space, enabling smooth gradient flow and efficient information transfer and complementarity across modalities, while also accelerating convergence. These results further validate the analysis in Sec. 3.4.
@@ 5.6 Different Batch Sizes $\Rightarrow$ Stability Analysis
In CLIP-based contrastive methods, different batch sizes lead to significant variations in the number of negative samples. As illustrated in Fig.3, batch size significantly affects the performance of various methods trained by feng2025anytouch dataset, evaluated across three subsets of TAG dataset. Notably, larger batch sizes tend to yield better performance, which aligns with observations from previous studies (Chen et al. 2020; Kerr et al. 2022). However, we also observe that when the batch size reaches 128, the performance of TLV-Link and VIT-LENS-2 no longer improves and may even decline. This is because a larger batch contains more data from different sensors, making it harder for single-sensor methods to handle the increased heterogeneity in tactile data. Overall, the proposed TLV-CoRe exhibits a smoother performance curve and demonstrates superior stability. This observation is consistent with the analysis presented in Sec. 3.4.
@@ 5.7 Ablation Study
We perform an ablation study on the components of TLV-CoRe trained by feng2025anytouch. As shown in Fig. 4, removing the SAM leads to a performance drop, and eliminating the decoupled loss $\mathcal{L}_{DL}$ causes an even greater decline, highlighting the crucial role of decoupled learning in handling multi-sensor data. Additionally, removing the UBA module results in the worst performance, underscoring its critical role in the overall architecture.
Additionally, we further investigate the impact of hyperparameters on TLV-CoRe trained by feng2025anytouch dataset: the decoupled coefficient $\lambda_{DL}$, the number of UBA $N_{UBA}$, and the dimension of UBA shared layer $d_{shared}$. As shown in Table 5, both $\lambda_{DL}$ and $d_{shared}$ involve a trade-off between performance and generalization: a low $\lambda_{DL}$ weakens the decoupled effect, while a high value suppresses contrastive learning; a small $d_{shared}$ leads to insufficient sharing, whereas a large one tends to cause overfitting. In contrast, increasing $N_{UBA}$ generally improves performance, which indicates that it’s necessary to fully promote collaboration and communication between different modalities.
@@
To investigate the impact of tactile images with similar styles but from different sensors on model performance, we conduct experiments using the GelSight and DIGIT data from feng2025anytouch dataset, as they exhibit stylistic similarity. Specifically, we randomly sample 5,000 instances from each dataset and additionally sample 2,500 instances from each to form a combined GelSight + DIGIT dataset. Due to the limited data size, we set the batch size to 8. Table 6 compares the performance of the multi-sensor methods AnyTouch and TLV-CoRe. Results show that AnyTouch suffers a performance drop when shifting from single-sensor data (GelSight or DIGIT) to the mixed dataset, indicating its inability to handle style-consistent sensor differences, which leads to degraded multi-sensor representations. In contrast, TLV-CoRe mitigates this issue through decoupled learning and demonstrates more stable performance.
@@ Appendix A Convergence Analysis
To validate our theoretical analysis of convergence properties in Theorem 3.1, we compare the training performance of TLV-CoRe against two state-of-the-art approaches: TLV-Link (Cheng et al. 2025) and AnyTouch (Feng et al. 2025). Fig. 5 shows the test loss trajectories over 12 epochs on the TAG dataset. The results provide strong empirical support for our theoretical claims. First, TLV-CoRe exhibits the fastest convergence rate among all methods, achieving a stable loss value by epoch 8, while other approaches continue to fluctuate. Second, unlike TLV-Link which shows instability after epoch 8 with increasing loss, TLV-CoRe maintains consistent performance. The combined effect of these properties enables TLV-CoRe to achieve both lower final test loss and more reliable optimization behavior, creating a foundation for robust downstream performance in this datasets and others.